OpenStudy (anonymous):

145 yards increase to 165 yards. What is the percentage increase?

jimthompson5910 (jim_thompson5910):

165 - 145 = ??

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

now divide that difference by the original amount

jimthompson5910 (jim_thompson5910):

20/145 = ??

OpenStudy (anonymous):

7.25

jimthompson5910 (jim_thompson5910):

divide the other way around

jimthompson5910 (jim_thompson5910):

the answer is NOT 7.25 %

OpenStudy (anonymous):

oh the other answer is 0.138

OpenStudy (anonymous):

13.79%?

jimthompson5910 (jim_thompson5910):

looks like they had a trap set up

jimthompson5910 (jim_thompson5910):

20/145 = 0.1379 = 13.79%

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

how does 0.1379 equal 13.79

jimthompson5910 (jim_thompson5910):

sorry you multiply 0.1379 by 100 to convert it to a percentage

jimthompson5910 (jim_thompson5910):

that's why 0.1379 = 13.79%

OpenStudy (anonymous):

oh ok thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

can you help me with some other math probs

jimthompson5910 (jim_thompson5910):

sure just a few though

OpenStudy (anonymous):

ok do you want me to put it in a new question box so you can get the medal

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

bike travels 12.5 miles in 2 hours and 30 minutes and 27.5 miles in 5 hours and 30 minutes. What is the constant of proportionality?

jimthompson5910 (jim_thompson5910):

2 hours, 30 minutes = 2.5 hours

jimthompson5910 (jim_thompson5910):

so you just divide 12.5 miles by 2.5 hours this will give you the speed, which is your answer

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok Ruth wants to buy a sweater. The sweater costs $39.50, and the sales tax in her state is 4%. In which of the following equations does x represent the amount of tax Ruth will pay?

jimthompson5910 (jim_thompson5910):

show me the equations and what you think

OpenStudy (anonymous):

ok wait a min

jimthompson5910 (jim_thompson5910):

k

jimthompson5910 (jim_thompson5910):

hints: part/whole = percent/100 and 4% = 4/100 = 1/25

OpenStudy (anonymous):

is it a

jimthompson5910 (jim_thompson5910):

what is x

jimthompson5910 (jim_thompson5910):

not the number but what does it represent

OpenStudy (anonymous):

the amount of tax Ruth will pay

jimthompson5910 (jim_thompson5910):

is that the "part" or the "whole"

OpenStudy (anonymous):

whole

OpenStudy (anonymous):

or is it part

jimthompson5910 (jim_thompson5910):

which one?

jimthompson5910 (jim_thompson5910):

think about it: the tax is a small piece of the picture (compared to the value of the item being bought)

OpenStudy (anonymous):

i dont know i hate percent problems

jimthompson5910 (jim_thompson5910):

the tax is the "part", the cost of the sweater is the "whole"

OpenStudy (anonymous):

so its the part

jimthompson5910 (jim_thompson5910):

the bigger number is always the "whole"

OpenStudy (anonymous):

its b the answer is b

jimthompson5910 (jim_thompson5910):

so part/whole = percent/100 part/whole = 4/100 ... because it's a 4% tax x/whole = 4/100 x/39.50 = 4/100 x/39.50 = 1/25

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok one more question

OpenStudy (anonymous):

The price of a large pizza in 2009 was $15, it was raised to $21 by 2011. By what percentage did the price of a large pizza increase?

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

what's the change in price

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

divide that by the original price

jimthompson5910 (jim_thompson5910):

you should get a number between 0 and 1

OpenStudy (anonymous):

0.4 then i multiply by hundred and that's my answer which is 40

jimthompson5910 (jim_thompson5910):

yep 40%

OpenStudy (anonymous):

thank you very much you are a great teacher

jimthompson5910 (jim_thompson5910):

glad to be of help